OpenStudy (anonymous):

The sum of 3 numbers is 79. The second number is 9 times the first and the third number is three more than the second. Find the numbers

hero (hero):

Hint: x = 1st number 9x = 2nd number 9x + 3 = 3rd number x + 9x + (9x + 3) = 79

hero (hero):

Solve for x

OpenStudy (anonymous):

Thanks!

hero (hero):

You're not done until you solve for x

hero (hero):

Do you know how to do that?

OpenStudy (anonymous):

I simplified and did \[10x +\left( 9x + 3 \right) = 79\]Now I'm kinda confused

OpenStudy (anonymous):

@Hero

hero (hero):

Hint: Remove the parentheses and continue solving.